OpenStudy (anonymous):

Solve the differential equation y prime =4x sqrt( 1-y^2).

myininaya (myininaya):

Try separating the variables then integrating.

OpenStudy (johnweldon1993):

Looks separable to me \[\large \frac{dy}{dx} = 4x\sqrt{1 - y^2}\] \[\large \frac{1}{\sqrt{1 - y^2}}dy = 4xdx\] Should be good from there!

OpenStudy (johnweldon1993):

Hint* Think of inverse trig functions whilst integrating

OpenStudy (anonymous):

would the 4x side be just 4? or would it be 2x^2 ?

OpenStudy (johnweldon1993):

You tell me \[\large \int 4xdx = 4\int xdx = ?\]

OpenStudy (anonymous):

arcsine(y) = 2x^2 ?

OpenStudy (johnweldon1993):

Perfect....but forgot the +c of course :D \[\large sin^{-1}(y) = 2x^2 + c\] So now to solve for 'y' we...?

OpenStudy (anonymous):

how do i "get rid" of the sin ^ (-1) ?

OpenStudy (johnweldon1993):

Hint* \(\large sin(sin^{-1}(x)) = x\) Well..really big hint lol

OpenStudy (anonymous):

y = sin(2x^2) + c

OpenStudy (johnweldon1993):

Uhh, just throw the constant inside of the parenthesis and you're good

OpenStudy (anonymous):

Explain why the initial value problem with y(0) = 4 does not have a solution

OpenStudy (johnweldon1993):

Is that a second part to this problem?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Okay so think about it...we have \[\large y = sin(2x^2 + c)\] If we have y(0) = 4 that means \[\large y(0) = sin(2(0)^2 + c)\] \[\large 4 = sin(0 + c)\] \[\large 4 = sin(c)\] Tell me a number to plug in for 'c' there that will result in an answer of 4 :)